# The Geometry of LLM Quantization: GPTQ as Babai’s Nearest Plane Algorithm

### v2

@@ B.1 Step 1
$=$ $\left(\sum_{k=1}^{j-1}\mathbf{e}_{j}^{\top}\bm{L}^{-1}\mathbf{e}_{k}\mathbf{e}_{k}^{\top}\bm{\Delta}^{(j-1)}\right)+\mathbf{e}_{j}^{\top}\bm{L}^{-1}\mathbf{e}_{j}\bm{0}+\left(\sum_{k=j+1}^{c}0\mathbf{e}_{k}^{\top}\bm{\Delta}^{(j-1)}\right)$ (Eqs. 1824)
@@
$=$ $\left(\sum_{k=1}^{j-1}\mathbf{e}_{j}^{\top}\bm{L}^{-1}\mathbf{e}_{k}\mathbf{e}_{k}^{\top}\bm{\Delta}^{(j)}\right)+\left(\sum_{k=j+1}^{c}\mathbf{e}_{j}^{\top}\bm{L}^{-1}\mathbf{e}_{k}\mathbf{e}_{k}^{\top}\bm{\Delta}^{(j)}\right)+\mathbf{e}_{j}^{\top}\bm{L}^{-1}\mathbf{e}_{j}\mathbf{e}_{j}^{\top}\bm{\Delta}^{(j)}-\mathbf{e}_{j}^{\top}\bm{\Delta}^{(j)}$ (Eqs. 1924)
@@ B.3 Step 3
$\tilde{\bm{\omega}}^{(j)}=$ $\mathbf{e}_{j}^{\top}\bm{D}_{\mathrm{U}}^{\frac{1}{2}}\bm{Y}^{(j+1)}$ (Eq. 34) (51)
$=$ $\mathbf{e}_{j}^{\top}\bm{D}_{\mathrm{U}}^{\frac{1}{2}}\bm{D}_{\mathrm{U}}^{-\frac{1}{2}}\bm{U}^{-1}\left(\tilde{\bm{W}}-\left(\sum_{k=j+1}^{c}\mathbf{e}_{k}\mathbf{e}_{k}^{\top}\right)\tilde{\bm{Q}}^{(j+1)}\right)$ (Eq. 50)
@@
$=$ $\mathbf{e}_{j}^{\top}\bm{U}^{-1}\tilde{\bm{W}}-\left(\left(\sum_{k=1}^{c}\mathbf{e}_{j}^{\top}\bm{U}^{-1}\mathbf{e}_{k}\mathbf{e}_{k}^{\top}\right)-\left(\sum_{k=1}^{j-1}0\mathbf{e}_{k}^{\top}\right)-1\mathbf{e}_{j}^{\top}\right)\tilde{\bm{Q}}^{(j+1)}$ (Eqs. 4345)
@@
$=$ $\mathbf{e}_{j}^{\top}\bm{U}^{-1}\tilde{\bm{W}}-\left(\sum_{k=1}^{c}\mathbf{e}_{j}^{\top}\bm{U}^{-1}\mathbf{e}_{k}\mathbf{e}_{k}^{\top}\right)\tilde{\bm{Q}}^{(j+1)}+\mathbf{e}_{j}^{\top}\tilde{\bm{W}}$ (Eq. 49)
@@
B.4 Proof of ineffectiveness of additional GPTQ refinement on Babai’s algorithm
@@
Appendix C Further Discussion on Quantization Error Bound
C.1 Proof of Absolute and Relative GPTQ Quantization Error Bounds
@@
For the relative no-clipping quantization error bound, we can plug in $\left\|\tilde{\bm{a}}_{(i)j}\right\|=\left|\bm{A}_{(i)}[j,j]\right|=\sqrt{\bm{D}_{(i)}[j,j]}=\sqrt{\left(\mathrm{diag}\left(\bm{T}^{-1}\bm{s}_{i}\right)\bm{D}\mathrm{\ diag}\left(\bm{T}^{-1}\bm{s}_{i}\right)\right)[j,j]}=\sqrt{\bm{D}[j,j]}\left|\left(\bm{T}^{-1}\bm{s}_{i}\right)[j]\right|:=d_{j}$ into Babai’s relative error bound in Section˜3.2.
C.2 Expected Quantization Error over a Uniform Hyper-Cuboid
@@
Expected residual norm. Using independence,
@@
C.3 Empirical Verification on Quantization Order and Error Bound
Changing the quantization order alters the diagonal matrix $\bm{D}$ of the LDL decomposition of the permuted Hessian and therefore the no-clipping GPTQ/Babai bound (see Section˜4.5). When per-group scales are approximately uniform, minimizing $\mathrm{tr}\left(\bm{D}\right)$ is a good proxy for tightening this bound. To assess different orders (back-to-front, front-to-back, random order, GPTQ’s act-order, and our min-pivot order), we run the calibration dataset from Section˜D.2 through the full-precision Qwen3-8B model and compute per-layer Hessians and calculate the $\mathrm{tr}\left(\bm{D}\right)$. For the random order, we average the results over 100 runs. Table˜2 reports $\mathrm{tr}\left(\bm{D}\right)$ for the layers in transformer block 18; other blocks and models show similar patterns. In block 18, act-order already reduces $\mathrm{tr}\left(\bm{D}\right)$ relative to the back-to-front/front-to-back/random baselines, especially in the Q·K·V and Gate·Up layers ($\approx$35-50% lower). Our min-pivot heuristic consistently attains the smallest trace. In practice, this tightens the theoretical layer-wise error bound and yields modest but consistent improvements. We can use act-order as a cheap option and reserve min-pivot for cases where a tighter bound is required.
@@ C.3 Empirical Verification on Quantization Order and Error Bound
Appendix D Further Applications and Experimental Results
D.1 Overflow-Tolerant Quantization Algorithms
@@ D.2 Experiment Setup
D.3 Accuracy Results
@@
Table˜5 shows additional zero-shot results on the Qwen3-8B model for RTN, GPTQ, HRTN, and HPTQ. Additional HPTQ results on other Qwen3 models are in Tables˜10, 6, 7, 8, 9 and 11.
@@
Method Avg Bitwidth Perplexity
@@
Model Avg Bitwidth Perplexity
@@
Method Avg Bits Wino MMLU HSwag PIQA SciQ TQA
@@
BF16 Baseline 16 68.11 73.02 74.90 77.80 95.7 36.35 54.50
HPTQ 4.125 67.17 72.28 74.84 77.42 95.6 35.01 53.36
3.125 66.93 70.96 73.18 77.53 95.4 36.11 54.73
2.125 59.19 52.99 63.86 72.52 86.8 31.09 49.01
GPTQ 4.125 68.82 71.76 74.22 77.58 95.3 36.35 54.55
3.125 68.35 65.80 70.80 75.46 75.46 36.11 55.21
2.125 52.25 34.25 39.32 57.83 57.83 28.40 46.91
HRTN 4.125 67.56 72.15 74.72 76.99 94.2 36.47 56.46
3.125 66.22 67.85 72.36 76.12 93.7 35.13 53.68
2.125 51.22 33.91 49.27 65.78 76.8 30.48 51.78
RTN 4.125 67.17 69.71 74.30 75.90 94.5 36.84 55.77
3.125 57.93 47.90 55.41 70.89 87.1 34.03 52.76
2.125 49.08 22.95 26.04 51.63 21.2 24.11 47.33
SSQR-1% 4.445 68.43 72.12 74.18 77.04 95.2 37.58 55.81
3.445 68.11 68.46 71.89 75.84 95.5 38.19 55.95
2.445 51.85 26.71 47.22 61.64 69.8 28.40 43.88
SSQR-2% 4.765 67.25 72.27 74.30 77.97 95.5 35.62 53.47
3.765 67.40 69.66 72.86 76.22 95.1 33.90 53.05
2.765 55.72 37.48 56.79 66.76 83.8 27.54 45.54
SSQR-3% 5.085 67.72 71.89 74.51 77.53 95.6 36.47 54.46
4.085 65.59 69.88 73.21 77.31 94.3 37.82 55.34
3.085 59.19 49.32 60.99 69.59 86.4 29.50 48.53
SSQR-4% 5.405 69.53 72.63 74.58 77.31 95.1 36.23 53.60
4.405 67.48 69.51 73.28 76.61 94.9 37.21 54.81
3.405 61.25 54.07 64.45 72.80 89.5 31.33 50.46
SSQR-5% 5.725 68.27 72.23 74.59 77.42 95.2 35.86 53.76
4.725 67.48 70.76 73.58 76.71 95.5 35.37 52.91
3.725 62.59 58.67 66.15 73.23 90.8 31.21 50.25
@@
Table 8: HellaSwag (%) zero-shot results for Qwen3 models quantized with HPTQ.
Avg Bitwidth 0.6B 1.7B 4B 8B 14B
16 47.30 60.40 68.46 74.90 78.82
4.125 45.70 59.29 67.63 74.84 78.88
3.125 40.70 56.01 66.28 73.18 77.73
2.125 28.77 39.40 52.13 63.86 70.96
@@
D.4 Technical Details and Performance of SSQR’s CUDA Kernel

### v4

@@ C.1 Step 1
$=$ $\left(\sum_{k=1}^{j-1}\mathbf{e}_{j}^{\top}\bm{L}^{-1}\mathbf{e}_{k}\mathbf{e}_{k}^{\top}\bm{\Delta}^{(j-1)}\right)+\mathbf{e}_{j}^{\top}\bm{L}^{-1}\mathbf{e}_{j}\bm{0}+\left(\sum_{k=j+1}^{c}0\mathbf{e}_{k}^{\top}\bm{\Delta}^{(j-1)}\right)$
                                 (Eqs. 1824)
@@
$=$ $\left(\sum_{k=1}^{j-1}\mathbf{e}_{j}^{\top}\bm{L}^{-1}\mathbf{e}_{k}\mathbf{e}_{k}^{\top}\bm{\Delta}^{(j)}\right)+\left(\sum_{k=j+1}^{c}\mathbf{e}_{j}^{\top}\bm{L}^{-1}\mathbf{e}_{k}\mathbf{e}_{k}^{\top}\bm{\Delta}^{(j)}\right)+\mathbf{e}_{j}^{\top}\bm{L}^{-1}\mathbf{e}_{j}\mathbf{e}_{j}^{\top}\bm{\Delta}^{(j)}-\mathbf{e}_{j}^{\top}\bm{\Delta}^{(j)}$
                                 (Eqs. 1924)
@@ C.3 Step 3
$\tilde{\bm{\omega}}^{(j)}=$ $\mathbf{e}_{j}^{\top}\bm{D}_{\mathrm{U}}^{\frac{1}{2}}\bm{Y}^{(j+1)}$ (51)
                                 (Eq. 34)
$=$ $\mathbf{e}_{j}^{\top}\bm{D}_{\mathrm{U}}^{\frac{1}{2}}\bm{D}_{\mathrm{U}}^{-\frac{1}{2}}\bm{U}^{-1}\left(\tilde{\bm{W}}-\left(\sum_{k=j+1}^{c}\mathbf{e}_{k}\mathbf{e}_{k}^{\top}\right)\tilde{\bm{Q}}^{(j+1)}\right)$
                                 (Eq. 50)
@@
$=$ $\mathbf{e}_{j}^{\top}\bm{U}^{-1}\tilde{\bm{W}}-\left(\left(\sum_{k=1}^{c}\mathbf{e}_{j}^{\top}\bm{U}^{-1}\mathbf{e}_{k}\mathbf{e}_{k}^{\top}\right)-\left(\sum_{k=1}^{j-1}0\mathbf{e}_{k}^{\top}\right)-1\mathbf{e}_{j}^{\top}\right)\tilde{\bm{Q}}^{(j+1)}$
                               (Eqs. 4345)
@@
$=$ $\mathbf{e}_{j}^{\top}\bm{U}^{-1}\tilde{\bm{W}}-\left(\sum_{k=1}^{c}\mathbf{e}_{j}^{\top}\bm{U}^{-1}\mathbf{e}_{k}\mathbf{e}_{k}^{\top}\right)\tilde{\bm{Q}}^{(j+1)}+\mathbf{e}_{j}^{\top}\tilde{\bm{W}}$
                                 (Eq. 49)
@@
C.4 Proof of Ineffectiveness of Additional GPTQ Refinement on Babai’s Algorithm
@@
Appendix D Further Discussion on Quantization Error Bound
D.1 Proof of Absolute and Relative GPTQ Quantization Error Bounds
@@
For the relative no-clipping quantization error bound, we can plug in $\left\|\tilde{\bm{a}}_{(i)j}\right\|=\left|\bm{A}_{(i)}[j,j]\right|=\sqrt{\bm{D}_{(i)}[j,j]}=\sqrt{\left(\mathrm{diag}\left(\bm{T}^{-1}\bm{s}_{i}\right)\bm{D}\mathrm{\ diag}\left(\bm{T}^{-1}\bm{s}_{i}\right)\right)[j,j]}=\sqrt{\bm{D}[j,j]}\left|\left(\bm{T}^{-1}\bm{s}_{i}\right)[j]\right|\coloneq d_{j}$ into Babai’s relative error bound in Section˜3.2.
D.2 Expected Quantization Error over a Uniform Hyper-Cuboid
@@
Expected residual norm. Using Lemma 7,
@@
D.3 Empirical Verification on Quantization Order and Error Bound
Changing the quantization order alters the diagonal matrix $\bm{D}$ of the LDL decomposition of the permuted Hessian and therefore the no-clipping GPTQ/Babai bound (see Section˜4.5). When per-group scales are approximately uniform, minimizing $\mathrm{tr}\left(\bm{D}\right)$ is a good proxy for tightening this bound. To assess different orders (back-to-front, front-to-back, random order, GPTQ’s act-order, and our min-pivot order), we run the calibration dataset from Section˜E.2 through the full-precision Qwen3-8B model and compute per-layer Hessians and calculate the $\mathrm{tr}\left(\bm{D}\right)$. For the random order, we average the results over 100 runs. Table˜2 reports $\mathrm{tr}\left(\bm{D}\right)$ for the layers in transformer block 18; other blocks and models show similar patterns. In block 18, act-order already reduces $\mathrm{tr}\left(\bm{D}\right)$ relative to the back-to-front/front-to-back/random baselines, especially in the Q·K·V and Gate·Up layers ($\approx$35-50% lower). Our min-pivot heuristic consistently attains the smallest trace. In practice, this tightens the theoretical layer-wise error bound and yields modest but consistent improvements. We can use act-order as a cheap option and reserve min-pivot for cases where a tighter bound is required.
@@ D.3 Empirical Verification on Quantization Order and Error Bound
Appendix E Further Applications and Experimental Results
E.1 Overflow-Tolerant Quantization Algorithms
@@ E.2 Experiment Setup
E.3 Accuracy Results for Qwen3 Models
@@
Table˜5 shows additional zero-shot results on the Qwen3-8B model for RTN, GPTQ, HRTN, and HPTQ. Additional HPTQ results on other Qwen3 models are in Tables˜9, 6, 7, 8 and 10.
@@
Method Avg. Bitwidth Perplexity
@@
Model Avg. Bitwidth Perplexity
@@
Method Avg. Bitwidth Wino MMLU PiQA SciQ TQA
@@
BF16 Baseline 16.000 68.11 73.02 77.80 95.7 36.35 54.50
HPTQ 4.125 67.17 72.28 77.42 95.6 35.01 53.36
3.125 66.93 70.96 77.53 95.4 36.11 54.73
2.125 59.19 52.99 72.52 86.8 31.09 49.01
GPTQ 4.125 68.82 71.76 77.58 95.3 36.35 54.55
3.125 68.35 65.80 75.46 75.46 36.11 55.21
2.125 52.25 34.25 57.83 57.83 28.40 46.91
HRTN 4.125 67.56 72.15 76.99 94.2 36.47 56.46
3.125 66.22 67.85 76.12 93.7 35.13 53.68
2.125 51.22 33.91 65.78 76.8 30.48 51.78
RTN 4.125 67.17 69.71 75.90 94.5 36.84 55.77
3.125 57.93 47.90 70.89 87.1 34.03 52.76
2.125 49.08 22.95 51.63 21.2 24.11 47.33
SSQR-1% 4.445 68.43 72.12 77.04 95.2 37.58 55.81
3.445 68.11 68.46 75.84 95.5 38.19 55.95
2.445 51.85 26.71 61.64 69.8 28.40 43.88
SSQR-2% 4.765 67.25 72.27 77.97 95.5 35.62 53.47
3.765 67.40 69.66 76.22 95.1 33.90 53.05
2.765 55.72 37.48 66.76 83.8 27.54 45.54
SSQR-3% 5.085 67.72 71.89 77.53 95.6 36.47 54.46
4.085 65.59 69.88 77.31 94.3 37.82 55.34
3.085 59.19 49.32 69.59 86.4 29.50 48.53
SSQR-4% 5.405 69.53 72.63 77.31 95.1 36.23 53.60
4.405 67.48 69.51 76.61 94.9 37.21 54.81
3.405 61.25 54.07 72.80 89.5 31.33 50.46
SSQR-5% 5.725 68.27 72.23 77.42 95.2 35.86 53.76
4.725 67.48 70.76 76.71 95.5 35.37 52.91
3.725 62.59 58.67 73.23 90.8 31.21 50.25
@@
E.4 Accuracy Results for Llama Models
Tables˜11, 12, 13, 14 and 15 report the evaluation results for Llama-3.2-3B-Instruct, Llama-3.1-8B-Instruct, and Llama-2-7B models under the same setups as in Section˜E.3.
Table 11: Perplexity of Llama-3.2-3B-Instruct model under HPTQ, GPTQ, and SSQR with different bitwidths.
Method Avg. Bitwidth Perplexity
WikiText-2 C4
BF16 Baseline 16.000 11.01 13.49
HPTQ 4.125 11.27 14.64
3.125 12.51 15.81
2.125 22.58 29.82
GPTQ 4.125 11.96 15.37
3.125 15.20 18.99
2.125 357.69 172.89
SSQR-1% 4.445 11.38 14.95
3.445 13.48 18.38
2.445 83.41 67.19
SSQR-2% 4.765 11.50 14.77
3.765 13.20 16.65
2.765 45.93 41.69
SSQR-3% 5.085 11.39 14.64
4.085 12.50 16.10
3.085 37.41 30.74
SSQR-4% 5.405 11.53 14.69
4.405 12.33 15.96
3.405 23.74 27.59
SSQR-5% 5.725 11.47 14.69
4.725 12.29 15.81
3.725 22.94 25.44
Table 12: Perplexity of Llama-3.1-8B-Instruct model under HPTQ, GPTQ, and SSQR with different bitwidths.
Method Avg. Bitwidth Perplexity
WikiText-2 C4
BF16 Baseline 16.000 7.20 9.09
HPTQ 4.125 7.37 9.99
3.125 7.84 11.04
2.125 11.89 16.37
GPTQ 4.125 7.56 10.46
3.125 9.44 13.16
2.125 148.15 71.33
SSQR-1% 4.445 7.50 10.30
3.445 8.67 12.35
2.445 57.26 39.96
SSQR-2% 4.765 7.48 10.20
3.765 8.32 11.75
2.765 25.18 25.21
SSQR-3% 5.085 7.41 10.11
4.085 8.16 11.54
3.085 17.27 20.03
SSQR-4% 5.405 7.39 10.05
4.405 8.01 11.31
3.405 13.22 17.77
SSQR-5% 5.725 7.38 10.03
4.725 7.98 11.13
3.725 12.12 16.13
Table 13: Perplexity of Llama-2-7B model under HPTQ, GPTQ, and SSQR-1% with different bitwidths.
Method Avg. Bitwidth Perplexity
WikiText-2 C4
FP16 Baseline 16.000 5.50 6.24
HPTQ 4.125 5.53 6.73
3.125 5.77 7.04
2.125 7.45 9.43
GPTQ 4.125 5.70 6.90
3.125 6.75 8.08
2.125 28.07 26.13
SSQR-1% 4.445 5.60 6.81
3.445 6.09 7.52
2.445 14.58 15.85
Table 14: Zero-shot evaluation results (%) for Llama-3.2-3B-Instruct under different quantization methods.
Method Avg. Bitwidth Wino MMLU PiQA SciQ HSwag
acc $\text{acc}_{\text{norm}}$
BF16 Baseline 16.000 68.75 62.18 76.17 95.4 53.27 71.65
HPTQ 4.125 68.03 61.57 76.55 95.0 53.02 71.28
3.125 68.35 58.50 7497 95.5 51.76 70.00
2.125 60.85 42.75 69.15 89.9 44.54 60.29
GPTQ 4.125 68.11 59.81 75.73 95.5 52.29 70.54
3.125 66.06 49.13 72.58 94.0 47.25 63.93
2.125 50.59 22.96 53.65 63.4 28.06 30.58
SSQR-1% 4.445 68.19 60.94 76.12 95.7 52.37 70.88
3.445 66.93 54.10 74.92 95.6 50.55 68.86
2.445 51.70 23.97 58.22 64.5 31.14 36.90
SSQR-2% 4.765 68.03 61.17 76.33 95.2 52.49 70.99
3.765 65.51 56.37 74.43 94.4 50.88 68.85
2.765 53.12 23.91 60.01 78.3 34.12 42.99
SSQR-3% 5.085 68.27 61.68 76.82 95.4 53.03 71.29
4.085 66.69 57.65 75.03 95.0 50.98 69.00
3.085 58.48 34.20 65.61 90.5 39.87 52.43
SSQR-4% 5.405 68.90 61.11 76.28 95.5 52.80 71.03
4.405 66.77 57.73 75.03 95.3 51.08 68.79
3.405 57.85 33.74 66.49 90.1 40.66 54.44
SSQR-5% 5.725 68.35 61.67 75.57 95.3 52.88 70.97
4.725 66.69 57.02 75.52 95.3 51.32 69.70
3.725 57.38 37.24 65.56 91.5 41.37 54.96
Table 15: Zero-shot evaluation results (%) for Llama-3.1-8B-Instruct under different quantization methods.
Method Avg. Bitwidth Wino MMLU PiQA SciQ HSwag
acc $\text{acc}_{\text{norm}}$
BF16 Baseline 16.000 73.72 68.31 80.14 97.3 59.81 79.59
HPTQ 4.125 73.56 67.90 79.49 97.7 59.57 79.25
3.125 72.77 64.58 79.16 96.9 58.42 78.21
2.125 63.69 45.01 69.15 90.8 49.84 67.98
GPTQ 4.125 73.80 65.68 79.27 97.2 58.61 78.36
3.125 72.45 58.19 77.37 95.5 55.21 74.57
2.125 54.93 24.67 54.46 75.1 31.77 37.79
SSQR-1% 4.445 74.43 66.78 79.65 96.9 59.18 78.93
3.445 72.45 60.14 77.97 96.3 56.74 76.24
2.445 52.80 23.07 58.49 74.1 33.25 40.05
SSQR-2% 4.765 73.80 67.21 79.49 97.2 58.94 78.53
3.765 73.24 63.13 78.78 96.4 57.63 77.22
2.765 54.30 27.08 61.04 82.5 38.41 50.41
SSQR-3% 5.085 72.93 67.38 79.54 96.9 59.64 79.07
4.085 73.09 63.77 79.11 96.6 57.62 77.40
3.085 54.54 26.15 58.81 83.6 38.34 49.52
SSQR-4% 5.405 73.24 66.95 79.92 96.9 59.32 79.06
4.405 73.24 62.92 78.73 96.5 57.61 77.47
3.405 54.54 29.95 54.95 82.3 39.80 51.87
SSQR-5% 5.725 74.03 67.91 80.52 97.2 59.49 79.39
4.725 73.40 64.14 79.05 97.0 58.16 77.63
3.725 64.25 42.59 72.58 88.7 49.94 68.20
E.5 Comparison with Other Quantization Methods
We compare zero-shot WinoGrande and PiQA accuracies of our methods (HPTQ, SSQR) against GPTQ and state-of-the-art post-training, weight-only quantizers AQLM (Egiazarian et al., 2024), QuIP# (Tseng et al., 2024a), and QTIP (Tseng et al., 2024b) on Llama-2-7B. Results are reported in Table˜16, sorted by average bitwidth. Metrics for AQLM, QuIP#, and QTIP are taken from their respective papers.
As shown in Table˜16, for average bitwidth $\geq$ 4, all methods yield accuracy close to the full-precision baseline. In the 3-4 bit regime, vanilla GPTQ falls behind recent methods; however, HPTQ and SSQR close this gap, bringing a scalar quantization approach to parity with vector quantization methods (AQLM, QuIP#, QTIP). In the 2-3 bit regime, HPTQ remains competitive with the state of the art.
Table 16: Comparing the zero-shot results of different quantization methods on Llama-2-7B.
Method Avg. Bitwidth WinoGrande PiQA
FP16 Baseline 16.000 69.46 78.13
AQLM 5.020 67.40 78.29
SSQR-1% 4.445 68.82 78.35
HPTQ 4.125 69.61 77.75
GPTQ 4.125 68.82 77.97
AQLM 4.040 67.32 78.24
QuIP# 4.000 67.60 78.40
QTIP 4.000 67.10 78.40
SSQR-1% 3.445 65.43 77.15
HPTQ 3.125 67.72 77.80
GPTQ 3.125 64.96 73.88
AQLM 3.040 66.93 76.88
QuIP# 3.000 66.50 77.30
QTIP 3.000 66.90 78.10
SSQR-1% 2.445 50.04 56.15
AQLM 2.290 65.67 74.92
HPTQ 2.125 65.82 73.56
GPTQ 2.125 49.64 56.20
AQLM 2.020 65.67 74.76
QuIP# 2.000 64.90 75.10
QTIP 2.000 64.70 75.90
E.6 Technical Details and Performance of SSQR’s CUDA Kernel